OpenStudy (anonymous):

simplify 3 3/4 divided by 2 1/2

OpenStudy (anonymous):

3 3/4 = 15/4 2 1/2 = 5/2 15/4 ÷ 5/2 = 15/4 x 2/5 =3/2

OpenStudy (anonymous):

THANKS'